Find an equation of the line passing through the given points. Use function notation to write the equation.

​(3​,−2​) and ​(5​,2​)


1
Expert's answer
2022-02-14T13:00:59-0500

The equation of line passing through two points "(x_1,y_1)\ \text{and}\ (x_2,y_2)" is:

"y-y_1=\frac{y_2-y_1}{x_2-x_1} (x-x_1)"

Here two points are ​(3​,−2​) and ​(5​,2​).

Let "(x_1,y_1)=(3,-2), (x_2,y_2)=(5,2)"

Thus, equation of line is

"y+2=\frac{2+2}{5-3} (x-3)\\
y+2=2x-6\\
y=2x-8\\
f(x)=2x-8"